# Improving Transformer Performance for French Clinical Notes Classification Using Mixture of Experts on a Limited Dataset

### v2

@@ Abstract
Clinical and Translational Impact Statement— This study highlights the potential of customized MoE-Transformers in enhancing clinical text classification, particularly for small-scale datasets like French clinical narratives. The MoE-Transformer’s ability to outperform several pre-trained BERT models marks a significant stride in applying advanced NLP techniques to clinical data. These findings pave the way for more accurate and efficient clinical text processing, potentially improving patient care and clinical research. The study underscores the importance of model selection and customization in achieving optimal performance for specific clinical applications, especially with limited data availability and within the constraints of hospital-based computational resources.
11footnotetext: This work was supported in part by the Natural Sciences and Engineering Research Council (NSERC), in part by the Institut de Valorisation des donnees de l’Universite de Montreal (IVADO), in part by the Fonds de la recherche en sante du Quebec (FRQS), and in part by the Fonds de recherche du Quebec-Nature et technologies (FRQNT). 22footnotetext: Thanh-Dung Le and Rita Noumeir are with the Biomedical Information Processing Lab, Ecole de Technologie Superieure, University of Quebec, Canada, and also with the Research Center at CHU Sainte-Justine, University of Montreal, Canada (Email: thanh-dung.le.1@ens.etsmtl.ca).33footnotetext: Philippe Jouvet is with the Research Center at CHU Sainte-Justine, University of Montreal, Canada.
@@ 2.1 French Clinical Data at CHUSJ
The clinical decision support system (CDSS) system in the CHU Sainte Justine (CHUSJ) hospital aims to improve the diagnosis and management of acute respiratory distress syndromes (ARDS) in real-time by automatically screening data from electronic medical records, chest X-rays, and other sources. Previous studies have found that the diagnosis of ARDS is often delayed or missed in many patients [23], emphasizing the need for more effective diagnostic tools. To diagnose ARDS, three main conditions must be detected: hypoxemia, chest X-ray infiltrates, and absence of cardiac failure [24]. The research team at CHUSJ has developed algorithms for detecting hypoxemia [25], analyzing chest X-rays [26, 27], and identifying the absence of cardiac failure. In addition, the team has performed extensive analyses of machine learning algorithms for detecting cardiac failure from clinical narratives using natural language processing [28, 29]. Implementing these algorithms could increase ARDS diagnosis rates and improve patient outcomes.
@@ Acknowledgment
This work was supported by a scholarship from the Fonds de recherche du Quebec-Nature et technologies (FRQNT) to Thanh-Dung Le, and grants from the Natural Sciences and Engineering Research Council (NSERC), the Institut de valorization des donnees (IVADO), and the Fonds de la recherche en sante du Quebec (FRQS).

### v3

@@ Abstract
Clinical and Translational Impact Statement— This study highlights the potential of customized MoE-Transformers in enhancing clinical text classification, particularly for small-scale datasets like French clinical narratives. The MoE-Transformer’s ability to outperform several pre-trained BERT models marks a stride in applying NLP techniques to clinical data and integrating into a Clinical Decision Support System in a Pediatric Intensive Care Unit. The study underscores the importance of model selection and customization in achieving optimal performance for specific clinical applications, especially with limited data availability and within the constraints of hospital-based computational resources.
11footnotetext: This work was supported in part by the Natural Sciences and Engineering Research Council (NSERC), in part by the Institut de Valorisation des donnees de l’Universite de Montreal (IVADO), in part by the Fonds de la recherche en sante du Quebec (FRQS), and in part by the Fonds de recherche du Quebec-Nature et technologies (FRQNT). 22footnotetext: Thanh-Dung Le is with the Biomedical Information Processing Lab, Ecole de Technologie Superieure, University of Quebec, Canada, and also is with the Interdisciplinary Centre for Security, Reliability, and Trust (SnT), University of Luxembourg, Luxembourg (Email: thanh-dung.le@uni.lu).33footnotetext: Rita Noumeir is with the Biomedical Information Processing Lab, Ecole de Technologie Superieure, University of Quebec, Canada.33footnotetext: Philippe Jouvet is with the Research Center at CHU Sainte-Justine, University of Montreal, Canada.
@@ 1 Introduction
The main contribution of this work is the introduction of a carefully curated, small-scale French clinical text dataset from CHU Sainte-Justine (CHUSJ) Hospital, which can serve as a valuable benchmark for resource-constrained NLP research in healthcare. Moreover, our findings demonstrate the feasibility of deploying the proposed MoE-Transformer in a practical clinical decision support system (CDSS), even when both the available data and computational resources are minimal.
@@ 2.1 French Clinical Data at CHUSJ
The CDSS system in the CHUSJ hospital aims to improve the diagnosis and management of acute respiratory distress syndromes (ARDS) in real-time by automatically screening data from electronic medical records, chest X-rays, and other sources. Previous studies have found that the diagnosis of ARDS is often delayed or missed in many patients [23], emphasizing the need for more effective diagnostic tools. To diagnose ARDS, three main conditions must be detected: hypoxemia, chest X-ray infiltrates, and absence of cardiac failure [24]. The research team at CHUSJ has developed algorithms for detecting hypoxemia [25], analyzing chest X-rays [26, 27], and identifying the absence of cardiac failure. In addition, the team has performed extensive analyses of machine learning algorithms for detecting cardiac failure from clinical narratives using natural language processing [28, 29]. Implementing these algorithms could increase ARDS diagnosis rates and improve patient outcomes.
@@ 5 Misclassification Interpretability
6 Limitations
While large language models (LLMs) have undoubtedly advanced numerous NLP tasks [58], their deployment in specialized clinical environments remains impractical. In our setting, stringent data privacy regulations prohibit transferring sensitive patient information to external servers or APIs, precluding fine-tuning and secure deployment of large off-the-shelf LLMs. Additionally, these models demand extensive computational resources—typically high-memory GPUs or multi-GPU clusters—far exceeding the capacity of our NVIDIA Quadro P620 (2 GB GPU memory). Furthermore, because LLMs are generally trained on broad-domain corpora, they often struggle to generalize to pediatric clinical narratives without significant retraining, amplifying data, and computational requirements. Given that our dataset comprises only 5,444 short clinical narratives, any attempt to fine-tune such large models would likely lead to overfitting and suboptimal performance. For example, a recent study [59] confirms that employing Low-Rank Adaptation during fine-tuning has proven ineffective in overcoming these limitations. Hence, a more focused and computationally efficient model, such as our MoE-Transformer, is a more viable alternative under these constraints.
Similarly, although state-of-the-art LLMs like Mistral [60] and LLaMA 3.21b [61] have demonstrated impressive capabilities in general NLP tasks, several fundamental limitations hinder their integration into our CDSS at CHUSJ. First, even the smallest LLMs typically contain no fewer than 5 billion parameters [62], making them more than 40 times larger than our domain-specific models, specifically trained on French biomedical clinical texts. This vast difference in scale implies that, even in zero-shot settings, general-purpose LLMs are prone to hallucinations and unreliable predictions when applied to pediatric clinical narratives [63]. Second, while these models are trained on massive datasets ranging from 825 billion to 6.3 trillion tokens, our dataset comprises only 580,000 tokens, further questioning their effectiveness without extensive domain-specific fine-tuning. Third, the computational demands of even the smallest LLMs (e.g., requiring a minimum of 1024-core NVIDIA Ampere architecture GPU with 32 tensor cores, 16G DRAM) are incompatible with our hardware capabilities. Finally, strict hospital privacy regulations preclude cloud-based inference, ruling out the possibility of relying on externally hosted models or resource-intensive fine-tuning methods. Despite their success in broader NLP applications, LLMs are not a feasible solution for our specialized clinical setting, thereby underscoring the advantages of our MoE-Transformer for local, real-time clinical text classification.
@@ 7 Conclusion
Future work will further explore enhancements to the evaluation methodology to gain deeper insights into the model’s performance and fairness in clinical text classification. Specifically, plans include incorporating class-wise performance metrics [64] to reveal detailed strengths and weaknesses across different clinical categories, thereby identifying potential biases and underperforming segments. Complementing this, a qualitative evaluation framework employing text perturbation experiments—such as synonym replacements, misspellings, and grammatical variations—will be developed to assess the robustness and generalization capabilities of the MoE-Transformer compared to large language model baselines. Additionally, drawing on insights from recent work on extracting social determinants of health from clinical texts [65], comparative analyses will be conducted to benchmark this approach against state-of-the-art methodologies. Recognizing the critical importance of fairness, future studies will also systematically examine potential biases in data curation, compilation, and model development by leveraging established frameworks like the AI Risk Repository, as highlighted in [66]. These explorations are expected to validate the approach’s practicality and robustness while contributing to developing fair, responsible, and practical AI applications in healthcare.
@@ Acknowledgment
This work was supported by a scholarship from the Fonds de recherche du Quebec-Nature et technologies (FRQNT) to Thanh-Dung Le, and grants from the Natural Sciences and Engineering Research Council (NSERC), the Institut de valorization des donnees (IVADO), and the Fonds de la recherche en sante du Quebec (FRQS). Data and reproducible codes are available upon request from Prof. Philippe Jouvet, M.D., PhD (Email: philippe.jouvet.med@ssss.gouv.qc.ca).